The value of a stock in dollars at the end of each month in 1992 varied according to the equation c(x) = 1/4x^2 - 2x + 6 where x represents the number of months.
a)For how many months did the stock decrease?
b)For how many months did the stock increase?

Answer by ewatrrr(24785)

 
Hi
The value of a stock in dollars at the end of each month in 1992 varied according to the equation
c(x) = 1/4x^2 - 2x + 6 = 1/4(x-4)^2 + 2 where x represents the number of months.
a)For how many months did the stock decrease? 4 mos
b)For how many months did the stock increase? 8mos
Might recommend Your downloading this FREE graph software

www.padowan.dk